Jefferson Lab > Physics > User Liaison
Privacy and Security Notice

Users Group: Board of Directors Minutes

Minutes of the meeting on November 11, 2004

BOD Members Present:
Paul Stoler (Chair), Gordon Cates (Chair Elect), David Armstrong, John Arrington Jian-Ping Chen, Ron Gilman, Cynthia Keppel Peter Monaghan, Clara Perdue, Julie Roche Christian Weiss (sitting in for Marc Vanderhaeghen)

Jefferson Lab Representatives Present:
Christoph Leemann, Swapan Chattopadhyay, Allison Lung, Larry Cardman, Dennis Skopik, Rusty Sprouse, Kees de Jager, Volker Burkert, Steve Wood, Hari Areti, Lia Merminga

The meeting began with an executive session at 8:30 AM. Minutes were taken by Arrington and Monaghan.

LAB STATUS Report: Christoph Leemann

Brief overview of status of experiments: DVCS experiments behind schedule but currently running well in Hall A, G11 experiment in Hall B is making progress in the analysis of their data, Hall C is running smoothly. At the moment, running at 5.75 GeV leads to a higher trip rate due to loss of overhead due to damage to cryomodules during the hurricane.

Recent accomplishments

JLab management contract open to competition soon. SURA is putting together a plan to satisfy the requirements of both the physics community and the DOE. Special focus on business processes and safety, especially in light of recent accidents at DOE labs. JLab has had a poorer than average safety record recently, and significant effort has been going into trying to improve the situation to help keep this from being a problem in the competition. Goal is to get from current rate of reportable incidents from 1.6 (per 200,000 person-hours) to 0.5.

12 GeV upgrade: Work continues toward CD-1. Management welcomes user involvement in the upgrade process. The goal is to put together the best possible case for the proposed physics program. Money is tight, so if we don't want to lose important physics, we need to explore possibilities that might help the situation: finding outside sources of funding (NSF or foreign contributions), coming up with ways to save money on the upgrade or equipment without compromising the physics program. The lab plans to campaign as aggressively as possible for full funding. If funding is such that we have to make difficult choices, the users will be involved in evaluating what physics will have to be lost, but it is premature to be trying to do this kind of planning now. The focus should be on building a strong, unified, case for the science that we want to do with the upgrade.

There were also discussions on the importance of keeping users interested in the 12 GeV upgrade, as well as finding new groups to work on developing the 12 GeV program. On the DOE side, the project does have momentum and the upgrade is "real". We need to make sure that it also has momentum in the physics community. At the same time, it is important to maintain a healthy and active 6 GeV program. Existing user groups may take up other projects and outside groups may not be attracted to JLab if we let the current program fall behind in anticipation of the upgrade.


Showed an updated version of the IPT (integrated project team). Still looking to fill some positions, and in particular would like suggestions from the users for someone to be the Physics Project Manager (the position held by John Domingo for the original CEBAF construction).

Discussed some R&D projects for 12 GeV. Development of the magnets and testing of superconducting cables/leads for Halls A and C, prototyping of detectors for Halls B and D.

Projected timeline for major reviews (carried out by JLab or DOE). Most of these times are not yet finalized.

Oct 22-24, 2004 - GlueX detector review (JLab+consultants)
Jan 2005 - Review of 'new' physics for CDR (JLab PAC)
Mar/Apr 2005 - 12 GeV science program (DOE)
Apr 2005 - Cryomodule design review (JLab)
May 2005 - Design/cost/schedule review (JLab)
July 2005 - "Lehman" review (DOE) - dry run in early may
Sep 2005 - Office of Science ESAAB review for CD-1 approval. Approve baseline range of project, full review of cost/schedule/performance.


Addition of 61,000 square feet of space in CEBAF center, and demolition of 22,000 square feet of trailer space (several of the stand-alone trailers and the portion of trailer city closest to CEBAF center (roughly half of trailer city). Timeline:

Furnishings for office space. Current plan is that most user space will be cubicles with two desks, but there will be some larger rooms with several desks and computers. A list of requirements for furnishings will be put together by December 2004, there will be mockup presentations by vendors in February 2005, and a contract will be awarded in April. Much discussion and a strong statement that it was considered extremely important to have doors that can be locked. Thia Keppel is the UGBOD representative on the committees that will make decisions on furnishings.

Discussion of other issues: More (different?) bike racks, especially at popular entrances. Discussions of how to determine who is moved into the new office space and how the move will be coordinated since at any given moment many users who have space in the trailers are not present at JLab.


Recent highlights.

Upcoming challenges.

  1. Energy: Can run at 5.75 GeV but with increased trip rate due to damage from the hurricane. Used to be able to achieve 5.75 with 'standard' trip rate (10/hour). Now limited to 5.5 GeV with normal trips, 5.75 with 15 trips/hr. July 2005 - Replace NC11 with 12 GeV prototype --> expect 5.75 GeV with 10 trips/hr July 2006 - Refurbishment of cryomodules --> expect 6.0 GeV with 10 trips/hr.
  2. Parity quality beam: Many of the specs for the PREX and Qweak experiments have been met, but requirements for position and beam charge asymmetries are somewhat better than achieved so far.
  3. Polarization: High polarization has been delivered, but the quantum efficiency has limited current. Want to improve stability and lifetime and make the superlattice cathode the standard.

Future: Want to move from a "construction" philosophy to an "operations" philosophy. In the past, people have focussed on building, maintaining, and optimizing individual subsystems. Want to move to a focus on optimizing the overall system to increase stability, help meet strict requirements of parity experiments.

Working on a new user web page. Detailed information on machine capabilities, help identify what experimental requirements will need either special care or additional development. New online forms for beam requirements of new proposals (to replace the existing forms).

Trying to build more expertise in beam diagnostics. Plans to grow to a three person group.

There was very positive feedback for accelerator from users, especially related to the high quality of parity beams. Discussion of what users can do to help. Close interactions between physics and accelerator have been very useful in developing parity quality beam for previous experiments. Also important to provide motivation for "service work" to keep people interested and involved. Some examples were publications for accelerator R&D work done collaboratively between the experiments and accelerator, R&D ideas from users, R&D money from user groups.

CDR Preparation/PAC27 ISSUES: Larry Cardman

PAC27 will have two parts: an update on the CDR physics case followed by a shortened PAC for reviewing new proposals. Current plan is to contact the editorial board (largely the same as the board for the pCDR) and have an updated executive summary that can be presented to the PAC members. At the meeting there will be presentations on each of the main topics in the CDR, with an emphasis on new physics and new experiments, updated presentations of the overall physics case, for final comment by the PAC before they are included in the CDR.


Next annual meeting: P. Stoler

Tentative theme/title for the workshop is "Challenges in QCD". Want to focus on the 'broader' physics topic, making sure to include speakers from outside of JLab, as well as speakers on the present and 12 GeV JLab program. Will try to have David Gross give an opening talk - In the short term, G. Cates will try to contact him and will look at setting a date for the workshop.

DNP Users' group meeting: R. Gilman

Good turnout at the DNP satellite meeting (estimated 30 people). Discussion focussed on the 12 GeV upgrade, with questions about the timetables, CDR preparation, and funding scenarios. Many questions about the procedures for decision making.

Computer issues: R. Gilman

Scientific computing meeting schedule for Nov 18th. Expect discussion of problems/praise for enterprise linux changeover. Some discussion of problems of with off-site access (slow response). There is a problem with too much interactive use of the ifarm machines, and there may be a change in policy to try and address this.

Graduate student issues: P. Monaghan

Discussion of space/office issues for students who are currently in trailers (also covered during status report on CEBAF center upgrade). The graduate student association (GSA) is planning to make several changes. First is to provide information and assistance to students (and postdocs) when they first move into the area. Discussing including postdocs in the GSA, organizing more activities than in the past. There are new members of the GSA who have agreed to help increase the role of the organization.

Quality of life: J. Arrington

Security issues led to the early removal of the computers from the residence facility last summer. The resfac wireless network was upgraded last summer, but there are still rooms where the reception is extremely weak (on the verge of useless). New thinclients running windows XP were installed a few days before the meeting. Early feedback was positive - much faster than the previous machines, a seamless X-client allowing one to run x applications from the JLab computers. It appears that not all of the software that is supposed to be installed is currently available. Looking for user feedback on the resfac machines.

Running experiments: J-P. Chen

The trip rate has been high for the present experiments running at 5.75 GeV (see accelerator status report above). All of the halls are currently running well, but Halls A and B lost time early on and are trying to get additional time to make up for this. Hall C is running ahead of schedule, largely due to the fact that the helium targets were able to take higher currents than expected.

There were some complaints that experiments and mcc sometimes fail to properly take restoration time into account after beam studies periods, delaying recovery from these down times. Hall B reported problems with saving and restoring machine configurations - Hari Areti responded that this was determined to be a communication issue between Hall B and MCC. There have been complaints about poor communication during accelerator down times - it was suggested that the operators make regular ELOG entries so that the halls can have a better feeling for what's going on during extended downtimes.

Finally, one last problem was brought up involving the need for better communication when problems involved three-way communication between the Halls, MCC, and the cryo group. There was a problem the previous week where cryo discovered what they thought was a problem in Hall C, and then contacted MCC about the problem and asked for a Hall C access. It took several phone calls and 20 minutes before anyone (MCC, the run coordinator, or the shift crew) could find out what the problem was and try to fix it.

Space issues: C. Keppel

There will be meeting in the near future to define the furniture requirements for the new office space in the CEBAF center addition, as well as discussions about how to allocate space and move people into the new offices when they become available.

Experimentalist/Theory Liaison: C. Weiss

No issues since the last meeting. Members of the theory group have been actively involved in the upgrade planning, both in putting together the physics case, writing the pCDR/CDR, and giving talks at JLab-wide and hall specific planning meetings.

PAC Issues: D. Armstrong

There were several complaints about the uncertainty related to PAC27 and the question of whether or not there would be review of new proposals. While it may be necessary or advantageous to skip regular PAC meetings, it was suggested that near the time of the PAC deadline, it should be made clear if there will be another PAC meeting in 12 months so that people have as much time as possible to decide if they have to submit in 6 months or if they can wait.

There was also a significant amount of discussion about the PAC decisions on pentaquark related proposals. Discussion of the PAC decision to wait on initial Hall B experiments before considering new pentaquark proposals and the fact that this reasoning was not included in the PAC25 report for experiments that were deferred, leading to some confusion on what the PAC thought were the reasons to defer some of the experiments. There as discussion of the role of management in giving guidance to the PAC on major programs. There was no resolution, and no plan to try and make any changes in the way the PAC evaluates proposals.

Post-Doc issues: J. Roche

No postdoc issues, but there are still visa issues for foreign users. In particular, there have been reports of disrespectful treatment of some users. Dennis Skopik says that he will investigate these reports.

Clara Perdue reports that things have improved over the last six months, and the problems in getting visas in a timely fashion have gotten better, although visitors from sensitive countries still have greater issues. Dennis Skopic reports that Clara will be leaving in December, and that there have been discussion on reorganizing User Liaison and visa issues - welcomes input on this issue.

HALL A REPORT: Kees de Jager

Happex-II and Happex-He both partially completed, limited to 30 uA due to septum heating. Discovered and removed foreign material in the He lines, and so expect to be able to reach higher currents for full run. With additional collimation, hope to achieve >65 uA.

DVCS on the proton has 85% of the requested data, after extending the run by six days. DVCS on the neutron is currently running well, but is short on time - will request to remove maintenance time after Thanksgiving.

Schedule for 2005 finalized: SRC with BigBite, installation of septa to finish hypernuclear running, installation of cryotargets to finish Happex-II and Happex-He. Removal of septa and running of GEn.

HALL B REPORT: Volker Burkert

g11 running complete (pentaquark experiment with hydrogen target). Tagger and detector calibration nearly complete. Primex installed and running - carbon running is done, lead running has started. eg3 scheduled to start in december. e1 (DVCS) is preparing to begin installation in february/march.

Long term schedule:
spring '05: DVCS run.
June '05: BONUS test run and g8 rungroup.
October '05: BONUS run if test goes well.
2006: frozen spin target, GDH at low Q, g12 experiment.

Status of pentaquark analysis (g10 and g11). Calibration is nearly competed for g11, and data cooking should begin soon (estimated 2 months to complete). g10 started 2 months earlier, finish first pass cooking of the data but found a tracking bug that induced a momentum bias in the results and so the data will have to be rerun. In both cased, there is roughly 7 TB of data, so a full replay takes something like 2-3 months.

HALL C REPORT: Steve Wood (for Rolf Ent)

Currently running a group of three (actually four) experiments: x>1, EMC effect in He, search for color transparency in pion production. These experiments are running ahead of schedule, largely because of higher than expected currents for Helium running, and so the subthreshold J/Psi measurement will take data during this run period.

Short and long term schedules dominated by large installation experiments (G0, GEp-III, SANE, Semi-SANE, Qweak, GEn), but will also try to fit in some of the measurements of R=sigma_L/sigma_T from the deuteron and nuclei (low-Q data is of particular interest to neutrino scattering community, so trying to get this part of the runs in early)

PAC27 has been given zero days to allocate in Hall C due to the large number of days approved at the previous PAC.